Can you help me solve this problem by using addition/elmination?
1.8x-2y=0.9
0.04x+0.18y=0.15

Answer by checkley77(12844)

1.8x-2y=0.9 multiply this equation by .18
0.04x+0.18y=0.15 multiply this equation by 2 & then add them.
.324x-.36y=.162
.08x+.36y=.30
-------------------
.404x=.462
x=.462/.404
x=.231/.202 ans.
Now multiply the second equation by -45 & add them again.
1.8x-2y=.9
-1.8x-8.1y=-6.75
------------------
-10.1y=-5.85
y=-5.85/-10.1
y=5.85/10.1 ans.
Proof:
.04*.231/.202+.18*5.85/10.1=.15
.00924/.202+1.053/10.1=.15
.0457+.1043=.15
.15=.15